Goods A and B are related goods.

The price of good A is $6.

When the price of good B is $29, demand for good A is P = 9 - 0.025QA.

When the price of good B is $31, demand for good A is P = 11 - 0.02QA.

Calculate the cross price elasticity of demand.
1
Expert's answer
2020-10-26T12:22:09-0400

cross elasticity of demand="\frac{percentage change in quantity demanded}{percentage change in price}"

% change in quantity demanded

"Q_1=9-(0.025\times 6)=8.85"


"Q_2=11-(0.02\times6)=10.88"


%change in Q=(10.88-8.85)"\times100" =203%


%change in price=(31-29)"\times100"=200%


Therefore; Cross price elasticity of demand= "\frac{203}{200}=1.015"